Trace the curve y^2 (2a - x ) = x^3
1
Expert's answer
2019-04-22T10:23:53-0400

1) Origin: Equation does not contain any constant term. Therefore, it passes through the origin.

The curve meets the coordinate axes only at the origin.

2) Symmetric about x-axis: The curve is symmetrical about x-axis, since only even powers of y occur.

3) Tangent at the origin: Equation of the tangent is obtained by equating to zero the lowest degree terms in the equation


"2ay^2-xy^2=x^3"

Equation of the tangent:


"2ay^2=0\Rightarrow y^2=0, y=0\ is\ a\ double\ point"

4) Cusp: As two tangents are coincident, therefore, origin is a cusp.

5) Asymptote parallel to y-axis: Equation of asymptote is obtained by equating the coefficient of highest degree of y to zero.


"y^2(2a-x)=x^3\Rightarrow 2a-x=0 \Rightarrow x=2a"

6) Region of absence of curve: "y^2" becomes negative on putting "x>2a" or "x<0," therefore, the curve does not exist for "x<0" and "x>2a."

We have cissoid.